# AnywhereDoor: Multi-Target Backdoor Attacks on Object Detection

### v1

@@ 5.3 Ablation Study
To evaluate our proposed techniques, we conduct experiments on Faster R-CNN and PASCAL VOC07+12 dataset with default hyperparameters illusrtated in Sec. 5.1.
As shown in Fig. 8, the radar chart compares different configurations on Clean mAP and five ASR metrics. The red polygon represents the model trained without objective disentanglement, achieving high Clean mAP but low ASR across all attack scenarios. If we remove trigger mosaicking, depicted by the green polygon, the model performs better in untargeted misclassification and generation but achieves only moderate ASR in other attack scenarios. In contrast, the model that discards strategic batching (depicted in yellow) improves ASR of all untargeted attacks but struggles with targeted scenarios, especially targeted misclassification, showing near-zero ASR.
The configuration that integrates all three techniques, AnywhereDoor, is shown in blue. Obviously, it achieves high ASR across all five attack scenarios while maintaining a decent Clean mAP. Its dominance over other configurations illustrates the effectiveness of our proposed techniques, resolving the limitations seen in complex attack scenarios, especially targeted misclassification, where it achieves $80\%$ higher ASR than baseline.
5.4 Parameter Study
Fig. 9 presents the performance of AnywhereDoor under different parameter settings. The experiments utilized the Faster R-CNN model and the PASCAL VOC07+12 dataset, with the default parameter settings described in Sec. 5.1. Overall, AnywhereDoor exhibits low sensitivity to parameter changes within a reasonable range, with performance metrics showing minor fluctuations within acceptable limits. It can be observed that the reasonable range for the learning rate is between $0.01$ and $0.15$, and the trigger generator is less sensitive to the learning rate compared to the object detection model. As the poisoning rate increases, various ASR metrics rise, while the Clean mAP slowly decreases. An epsilon value of at least $0.05$ is required to achieve satisfactory performance; further increases in epsilon do not significantly enhance ASR and instead make the trigger on the images visible, reducing the attack’s stealthiness. When epsilon is $0.05$, the trigger on the images is almost imperceptible to the naked eye, as illustrated by the visualization results in Tab. 5.
Figure 9: Impact of parameters on AnywhereDoor. (a) Learning rate of trigger generator network. (b) Poisoning rate, i.e., the fraction of backdoor samples in a mini-batch. (c) Epsilon, i.e., the maximum pixel change to clean image. (d) Trigger size, i.e., the side length of each trigger. Similar to Fig. 8, although both Clean mAP and ASR range from 0 to 100, they differ in meaning.
@@ 5.5 Resilience Against Defenses
Tab. 4 evaluates AnywhereDoor’s robustness against six common defenses, including input- and model-based mitigation methods. We use Faster R-CNN and PASCAL VOC07+12 with default hyperparameters from Sec. 5.1. JPEG compression quality is set to $85$. Both mean and median filters use a kernel size of $3$. Fine-tuning and fine-pruning undergo $12$ retraining epochs, with pruning-based methods applied at a pruning rate of $0.3$.
Mitigation-based Defense Methods Clean mAP Untar. Removal ASR Tar. Removal ASR Untar. Miscls. ASR Tar. Miscls. ASR Untar. Gen. ASR
No Defense 76.3 97.5 86.2 97.8 80.6 88.8
(a) Input JPEG Compr. 76.1 97.7 87.4 98.2 79.6 89.4
Mean Filter 75.3 99.7 91.0 98.4 81.0 89.2
Median Filter 75.8 99.7 92.5 98.8 82.2 89.8
(b) Model Fine-tuning 76.9 95.5 76.5 89.3 69.6 82.8
Pruning 26.8 61.1 61.8 90.0 9.1 13.1
Fine-pruning 26.8 61.1 62.2 90.0 9.0 13.1
Table 4: Evaluation of effectiveness of multiple mitigation-based defenses on AnywhereDoor.
Input-based Mitigation Defenses. Input sanitization methods, such as JPEG compression [15], mean filter, and median filter [58], aim to suppress triggers by modifying the input. As shown in Tab. 4 (a), these methods preserve clean mAP well but fail to eliminate the backdoor, as shown by the ASR values that remain high or even increase in some cases. This insenstivity of ASR indicates that AnywhereDoor’s trigger patterns are resilient to minor input distortions, making these defenses ineffective.
Model-based Mitigation Defenses. Model sanitization techniques, including Fine-tuning, Pruning, and Fine-pruning [39], were alternative approaches to counter backdoor attacks. Tab. 4 (b) shows that fine-tuning slightly reduces ASR in targeted misclassification but leaves most ASR values above $80\%$. Pruning and fine-pruning significantly lower ASR but also drastically reduce clean mAP to $26.8$, compromising model utility. This trade-off highlights that model-based defenses partially counteract backdoors but at a high cost to performance.
Overall, none of the these defense methods could completely eliminate the backdoor implanted by AnywhereDoor, underscoring the resilience of our approach to these mitigations.
@@ 6 Conclusion
We presented AnywhereDoor, a backdoor attack that allows multi-target manipulation of object detection models, introducing unprecedented flexibility beyond prior attacks reliant on predefined triggers and behaviors. By leveraging joint training of a trigger generator with the victim model and dynamic sample poisoning, AnywhereDoor overcomes challenges specific to object detection via three techniques: objective disentanglement, trigger mosaicking, and strategic batching. These address complexities including class-behavior associations, localized region extraction, and inter-class imbalance in datasets. Extensive experiments demonstrate AnywhereDoor’s effectiveness, achieving high ASR while preserving clean model performance, underscoring vulnerabilities in current object detection systems and the threats posed by backdoor attacks.
Appendix A Attack Success Rate Calculation
@@
The untargeted removal scenario (Alg. 1) focuses on eliminating bounding boxes from the clean predictions, accumulating the total number of objects ($t_{i}$) and the count of successfully removed ones ($s_{i}$). In the targeted removal scenario (Alg. 2), an additional parameter, the victim class $C_{v}$, is provided to specify which class is to be removed. Here, $t_{i}$ includes only the bounding boxes of the victim class in the clean predictions, and $s_{i}$ represents those successfully removed. An IoU threshold of $0.5$ is used to identify if two bounding boxes refer to the same object. For untargeted misclassification (Alg. 3), the aim is to alter the class labels of bounding boxes. We accumulate $t_{i}$ as the total number of bounding boxes in the clean predictions and $s_{i}$ as the count of bounding boxes whose classes are successfully changed. A successful misclassification occurs when a bounding box in the dirty prediction has a different class from the corresponding one in the clean prediction, provided it is not derived from another misclassified bounding box in the clean predictions. In targeted misclassification (Alg. 4), the attack targets a specific victim class $C_{v}$, aiming to misclassify its instances as the target class $C_{t}$. The total count $t_{i}$ is the number of victim-class bounding boxes in the clean predictions, while $s_{i}$ counts those successfully changed to $C_{t}$ in the dirty predictions. Success is determined by the presence of a corresponding bounding box with class $C_{t}$ in the dirty predictions and an IoU exceeding $0.5$. Finally, the ASR of untargeted generation (Alg. 5) is calculated in a sample-based way, evaluating whether new bounding boxes are generated . Each sample contributes $t_{i}=1$, and $s_{i}=1$ if the number of bounding boxes in the dirty predictions exceeds those in the clean predictions.
@@ 9       $S,T\leftarrow S+s_{i},T+t_{i}$ ;
Appendix B Transferability Study
We prepared pretrained trigger generators that were initially trained jointly with Faster R-CNN, DETR, and YOLOv3 on the PASCAL VOC07+12 dataset. These trigger generators were then paired with different models for another training process in which the parameters of trigger generators are frozen, with only models being trained. As shown in Fig. 10, subfigures (a), (b), and (c) display results for Faster R-CNN, DETR, and YOLOv3, respectively, each trained with a trigger generator originally paired with the same model, as well as with those from two other models. Each bar chart reports performance metrics of Clean mAP and ASR of five attack scenarios. Overall, both Clean mAP and ASR show minimal degradation when using different trigger generators, indicating that the attacks remain effective across model combinations. Notable ASR declines are observed primarily in the two targeted attack scenarios, highlighting their higher difficulty and lower ASR stability. The combination of Faster R-CNN and DETR maintains consistent performance, while YOLOv3 experiences more significant ASR drops when paired with the other two models. This suggests structural differences in YOLOv3 that impact its backdoor attack performance, aligning with the results presented in  Sec. 5.2 of the main paper.
@@ Appendix C Visual Samples
We provide additional visual examples to illustrate the effectiveness of AnywhereDoor across various attack scenarios. For the convenience of viewing, we select some examples with small number of classes and objects and large object size from both PASCAL VOC07+12 and MSCOCO, and use Faster R-CNN trained on PASCAL VOC07+12 to obtain visualization results on five attack scenarios, as shown in Tab. 5.

### v2

@@ 5.3 Ablation Study
To assess the importance of each component in AnywhereDoor, we conduct an ablation study by removing one component at a time. The radar chart in Figure 9 presents results for clean mAP and ASR across different attack scenarios. First, removing objective disentanglement (red) causes a significant drop in ASR across all scenarios, highlighting its role in scaling the number of attack targets that can be manipulated. Second, removing trigger mosaicking (green) has a substantial impact on targeted attacks, demonstrating the importance of maintaining trigger effectiveness across different subregions of the image. Third, removing strategic batching also causes the attack to fail in targeted scenarios, underscoring the importance of balanced learning opportunities during training. Based on this study, all components are essential to the overall effectiveness of AnywhereDoor.
@@ 5.4 Resilience Against Defenses
While defense resilience is not the primary objective of this paper, we surprisingly found that AnywhereDoor is resilient against some defenses. Table 4 evaluates AnywhereDoor’s robustness against seven methods, including input-based and model-based mitigation.
Mitigation-based Defense Methods Clean mAP Attack Success Rate
Removal Miscls. Gen.
Untar. Tar. Untar. Tar. Untar.
No Defense 53.1 97.5 86.2 97.8 80.6 88.8
(a) JPEG Compr. 53.0 97.7 87.4 98.2 79.6 89.4
Mean Filter 52.5 99.7 91.0 98.4 81.0 89.2
Median Filter 52.8 99.7 92.5 98.8 82.2 89.8
NEO 52.6 97.5 85.6 97.8 79.1 88.8
(b) Fine-tuning 53.6 95.5 76.5 89.3 69.6 82.8
Pruning 18.3 61.1 61.8 90.0 9.1 13.1
Fine-pruning 18.6 61.1 62.2 90.0 9.0 13.1
Table 4: While not being specifically designed for, AnywhereDoor demonstrates resilience against some defenses.
Input-based Mitigation. Input sanitization methods, such as JPEG compression [12], mean filter, median filter [46], and NEO [44], aim to suppress triggers by modifying the input. As shown in Table 4a, these methods preserve clean mAP well but fail to eliminate the backdoor, as indicated by the ASR, which remains high or even gets increased in some cases. This suggests that AnywhereDoor’s trigger patterns are insensitive to minor input distortions, making these defenses ineffective.
Model-based Mitigation. Model sanitization techniques, including fine-tuning, pruning, and fine-pruning [32], are alternative approaches to counter backdoor attacks. Table 4b shows that fine-tuning reduces ASR in targeted misclassification but leaves ASR in most other scenarios above $82.8\%$. Pruning and fine-pruning significantly lower ASR but also drastically reduce clean mAP to $18.3$, compromising model utility. This trade-off highlights that model-based defenses partially counteract backdoors but at a high cost to performance.
@@ 6 Conclusion
We presented AnywhereDoor, a backdoor attack that enables multi-target manipulation of object detection models, offering unprecedented flexibility beyond prior attacks that rely on a small number of predefined triggers and behaviors. By jointly training a trigger generator with the victim model and employing dynamic sample poisoning, AnywhereDoor overcomes challenges specific to object detection through three key techniques: objective disentanglement, trigger mosaicking, and strategic batching. Extensive experiments demonstrate AnywhereDoor’s effectiveness, achieving high ASR while preserving clean model performance and highlighting vulnerabilities in current object detection systems and the growing threat posed by backdoor attacks.
Supplementary Material
We provide the source code of AnywhereDoor at https://github.com/HKU-TASR/AnywhereDoor. This document provides additional details to support our main paper.
Appendix A Transferability Study
We prepared pretrained trigger generators that were initially trained jointly with Faster R-CNN, DETR, and YOLOv3 on the PASCAL VOC07+12 dataset. These trigger generators were then paired with different models for another training process in which the parameters of trigger generators are frozen, with only models being trained. As shown in Figure 10, subfigures (a), (b), and (c) display results for Faster R-CNN, DETR, and YOLOv3, respectively, each trained with a trigger generator originally paired with the same model, as well as with those from two other models. Each bar chart reports performance metrics of Clean mAP and ASR of five attack scenarios. Overall, both Clean mAP and ASR show minimal degradation when using different trigger generators, indicating that the attacks remain effective across model combinations. Notable ASR declines are observed primarily in the two targeted attack scenarios, highlighting their higher difficulty and lower ASR stability. The combination of Faster R-CNN and DETR maintains consistent performance, while YOLOv3 experiences more significant ASR drops when paired with the other two models. This suggests structural differences in YOLOv3 that impact its backdoor attack performance, aligning with the results presented in Section 5.2 of the main paper.
Appendix B Attack Success Rate Calculation
@@
The untargeted removal scenario (Algorithm 1) focuses on eliminating bounding boxes from the clean predictions, accumulating the total number of objects ($t_{i}$) and the count of successfully removed ones ($s_{i}$). In the targeted removal scenario (Algorithm 2), an additional parameter, the victim class $C_{v}$, is provided to specify which class is to be removed. Here, $t_{i}$ includes only the bounding boxes of the victim class in the clean predictions, and $s_{i}$ represents those successfully removed. An IoU threshold of $0.5$ is used to identify if two bounding boxes refer to the same object. For untargeted misclassification (Algorithm 3), the aim is to alter the class labels of bounding boxes. We accumulate $t_{i}$ as the total number of bounding boxes in the clean predictions and $s_{i}$ as the count of bounding boxes whose classes are successfully changed. A successful misclassification occurs when a bounding box in the dirty prediction has a different class from the corresponding one in the clean prediction, provided it is not derived from another misclassified bounding box in the clean predictions. In targeted misclassification (Algorithm 4), the attack targets a specific victim class $C_{v}$, aiming to misclassify its instances as the target class $C_{t}$. The total count $t_{i}$ is the number of victim-class bounding boxes in the clean predictions, while $s_{i}$ counts those successfully changed to $C_{t}$ in the dirty predictions. Success is determined by the presence of a corresponding bounding box with class $C_{t}$ in the dirty predictions and an IoU exceeding $0.5$. Finally, the ASR of untargeted generation (Algorithm 5) is calculated in a sample-based way, evaluating whether new bounding boxes are generated . Each sample contributes $t_{i}=1$, and $s_{i}=1$ if the number of bounding boxes in the dirty predictions exceeds those in the clean predictions.
@@ Appendix C Visual Samples
We provide additional visual examples to illustrate the effectiveness of AnywhereDoor across various attack scenarios. For the convenience of viewing, we select some examples with small number of classes and objects and large object size from both PASCAL VOC07+12 and MSCOCO, and use Faster R-CNN trained on PASCAL VOC07+12 to obtain visualization results on five attack scenarios, as shown in Table 5.
@@ Appendix D Limitations and Future Works
Appendix E Experimental Setup
Datasets
The experimental configuration employs two standard object detection benchmarks: PASCAL VOC07+12 and MSCOCO. For PASCAL VOC07+12, we combine the VOC2007 train set (2,501 images) with the VOC2012 trainval set (11,540 images) to form the training data, while using the VOC2007 val set (2,510 images) for testing. The MSCOCO dataset utilizes its full training set of 118k images (80 classes) for model training, with evaluation conducted on the validation set containing 5k images. For targeted misclassification attack, we select a 5-class subset consisting of person, car, bus, bicycle, and motorbike (equivalent to motorcycle in MSCOCO annotations) from both datasets.
Implementation Details
We utilize models pretrained on MSCOCO provided by the mmdetection [3]. The pretraining epochs are set as 12 for Faster R-CNN, 150 for DETR, and 273 for YOLOv3. The optimization framework employs distinct configurations for each architecture. Faster R-CNN utilizes SGD with learning rate 0.02, momentum 0.9, and weight decay 0.0001. DETR adopts AdamW optimizer ($\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999) with a lower learning rate 0.0001 and equivalent weight decay regularization (0.0001). YOLOv3 implements an SGD optimizer with learning rate 0.0001 and weight decay 0.0005, combined with a cosine learning rate scheduler to accommodate its multi-scale detection mechanism. The trigger generator for all models is trained with Adam and has a learning rate of 0.1.
We use input resolutions of $(1000,600)$ with batch size 8 for VOC and $(500,300)$ with batch size 16 for COCO, maintaining a fixed poisoning rate $p=0.5$ (4 samples per VOC batch and 8 per COCO batch). The adversarial trigger is implemented as a $3\times 30\times 30$ RGB patch with $\ell_{\infty}$-norm perturbation constraint $\epsilon=0.05$, generated by a 1-layer fully connected network optimized via Adam ($\beta_{1}=0.5$, $\beta_{2}=0.999$).